# Anosov representations, strongly convex cocompact groups and weak eigenvalue gaps

### v3

@@ 10. Examples and counterexamples
In this section, we discuss examples of representations of surface groups enjoyng some of the properties of Anosov representations which are not $P_{1}$-Anosov. The examples show that the assumptions of the main results of this paper are necessary. Throughout this section $S_{g}$ denotes a closed orientable surface of genus $g\geq 2$.
@@
$\bullet$ $\rho$ is a quasi-isometric embedding, $P_{1}$-divergent and preserves a properly convex domain $\Omega$ of $\mathbb{P}(\mathbb{R}^{12})$.
$\bullet$ $\rho$ admits continuous, injective, $\rho$-equivariant maps $(\xi_{1},\xi_{11}):\partial_{\infty}\pi_{1}(S)\rightarrow\mathbb{P}(\mathbb{R}^{12})\times\mathsf{Gr}_{11}(\mathbb{R}^{12})$ satisfying the Cartan property. The proximal limit set of $\rho(\pi_{1}(S_{g}))$ in $\mathbb{P}(\mathbb{R}^{12})$ is $\xi_{1}(\partial_{\infty}\pi_{1}(S))$ and does not contain projective line segments.
$\bullet$ $\rho$ admits continuous, $\rho$-equivariant transverse maps $(\xi_{4},\xi_{8}):\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathsf{Gr}_{4}(\mathbb{R}^{12})\times\mathsf{Gr}_{8}(\mathbb{R}^{12})$.
$\bullet$ $\rho$ is not $P_{k}$-Anosov for any $k=1,\ldots,11$.
The previous example shows that the assumption of transversality in Theorem 1.1 is necessary. Moreover, the maps $\xi_{4}$ and $\xi_{8}$ are transverse although $\rho$ is not $P_{4}$-Anosov, therefore Zariski density is also necessary in Theorem 1.3.
@@
Let $S_{g}$ be a closed orientable surface of genus at least $2$ and $\phi:S_{g}\rightarrow S_{g}$ a pseudo-Anosov homeomorphism. The mapping torus $M_{\phi}$ of $S_{g}$ with respect to $\phi$ is a closed $3$-manifold whose fundamental group is isomorphic to the HNN extension
$\pi_{1}(M_{\phi})=\Big{\langle}\pi_{1}(S_{g}),t\ \Big{|}\ tat^{-1}=\phi_{\ast}(a),\ a\in\pi_{1}(S_{g})\Big{\rangle}$
@@
where $\phi_{\ast}$ is the automorphism of $\pi_{1}(S_{g})$ induced by $\phi$. Thurston in [37] (see also Otal [35]) proved that there exists a convex cocompact reprsentation $\rho_{0}:\pi_{1}(M_{\phi})\rightarrow\mathsf{PO}(3,1)$. The representation $\rho_{0}$ lifts to a $P_{1}$-Anosov representation in $\mathsf{SL}_{4}(\mathbb{R})$ which we continue to denote by $\rho_{0}$ and let $\rho_{\textup{Fiber}}:=\rho_{0}|_{\pi_{1}(S_{g})}$. By a result of Cannon-Thurston [13], there exists a continuous equivariant surjection $\theta:\partial_{\infty}\pi_{1}(S_{g})\twoheadrightarrow\partial_{\infty}\pi_{1}(M_{\phi})$. By precomposing $\theta$ with the Anosov limit map of $\rho_{0}$ in $\mathbb{P}(\mathbb{R}^{4})$, we obtain a $\rho_{\textup{Fiber}}$-equivariant continuous map $\xi_{\textup{Fiber}}:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathbb{P}(\mathbb{R}^{4})$. Let $\gamma\in\pi_{1}(S_{g})$ be an element representing a separating simple closed curve on $S_{g}$. We may choose a Zariski dense, Hitchin representation $\rho_{\textup{H}}:\pi_{1}(S_{g})\rightarrow\mathsf{SL}_{3}(\mathbb{R})$ with $2\log\ell_{1}(\rho_{\textup{Fiber}}(\gamma))=\log\ell_{1}(\rho_{\textup{H}}(\gamma))$. We claim that $\rho=\rho_{\textup{Fiber}}\otimes\rho_{\textup{H}}:\pi_{1}(S)\rightarrow\mathsf{SL}_{12}(\mathbb{R})$ satisfies the required properties.
@@
Let $\otimes:\mathsf{SO}(3,1)\times\mathsf{SL}_{3}(\mathbb{R})\rightarrow\mathsf{SL}_{12}(\mathbb{R})$ be the irreducible tensor product representation $(g_{1},g_{2})\mapsto g_{1}\otimes g_{2}$. Let $G$ be the Zariski closure of $\rho_{\textup{Fiber}}\times\rho_{\textup{H}}$ into $\mathsf{SO}(3,1)\times\mathsf{SL}_{3}(\mathbb{R})$. Note that the projection of the identity component $G^{0}$ into $\mathsf{SO}(3,1)$ (resp. $\mathsf{SL}_{3}(\mathbb{R})$) is normalized by $\rho_{\textup{Fiber}}(\pi_{1}(S))$ (resp. $\rho_{\textup{H}}(\pi_{1}(S_{g}))$), so it has to be surjective. Since the Zariski closures of $\rho_{\textup{Fiber}}$ and $\rho_{\textup{H}}$ are simple and not locally isomorphic, it follows by Goursat’s lemma that $G=\mathsf{SO}(3,1)\times\mathsf{SL}_{3}(\mathbb{R})$. We conclude that $\rho$ is strongly irreducible.
@@
We obtain a properly convex domain $\Omega$ of $\mathbb{P}(\mathbb{R}^{12})$ preserved by $\rho(\pi_{1}(S_{g}))$ as follows. Let $\Omega_{1}$ and $\Omega_{2}$ be properly convex domains of $\mathbb{P}(\mathbb{R}^{4})$ and $\mathbb{P}(\mathbb{R}^{3})$ preserved by $\rho_{\textup{Fiber}}(\pi_{1}(S))$ and $\rho_{\textup{H}}(\pi_{1}(S_{g}))$ respectively, and $\Omega_{i}^{\prime}$ a properly convex cone lifting $\Omega_{i}$ for $i=1,2$. The compact set
$\mathcal{C}=\big{\{}[u_{1}\otimes u_{2}]:u_{i}\in\overline{\Omega_{1}^{\prime}},u_{i}\in\overline{\Omega_{2}^{\prime}}\big{\}}$
@@
is connected, spans $\mathbb{R}^{12}$ and is contained in an affine chart $A\subset\mathbb{P}(\mathbb{R}^{12})$. We finally take $\Omega$ to be the interior of the convex hull of $\mathcal{C}$ in $A$.
The representations $\rho_{\textup{Fiber}}$ and $\rho_{\textup{H}}$ are $P_{1}$-divergent hence $\rho$ is also $P_{1}$-divergent. Note that
$\sigma_{1}(\rho(\gamma))=\sigma_{1}(\rho_{\textup{Fiber}}(\gamma))\sigma_{1}(\rho_{\textup{H}}(\gamma))\ \ \forall\gamma\in\Gamma,$
@@
hence $\rho$ is a quasi-isometric embedding. Let $\xi_{\textup{H}}:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathbb{P}(\mathbb{R}^{3})$ and $\xi_{\textup{H}}^{-}:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathsf{Gr}_{2}(\mathbb{R}^{3})$ be the Anosov limit maps of $\rho_{\textup{H}}$. The map $\xi_{1}:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathbb{P}(\mathbb{R}^{12})$ defined as
$\xi_{1}(x)=[k_{x}e_{1}\otimes k_{x}^{\prime}e_{1}],$
@@
where $\xi_{\textup{Fiber}}(x)=[k_{x}e_{1}]$ and $\xi_{\textup{H}}(x)=[k_{x}^{\prime}e_{1}]$, is continuous and $\rho$-equivariant. Since $\rho$ is strongly irreducible, the proof of Corollary 4.6 shows that the map $\xi_{1}$ satisfies the Cartan property. The image of $\xi_{1}$ is the $P_{1}$-proximal limit set of $\rho(\pi_{1}(S_{g}))$ in $\mathbb{P}(\mathbb{R}^{12})$. Similarly, the dual reprsentation $\rho^{\ast}=\rho_{\textup{Fiber}}^{\ast}\otimes\rho_{\textup{H}}^{\ast}$ admits a $\rho^{\ast}$-equivariant map $\xi_{1}^{\ast}:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathbb{P}(\mathbb{R}^{12})$, so we obtain the $\rho$-equivariant map $\xi_{11}$.
@@
$\xi_{4}(x)=\mathbb{R}^{4}\otimes_{\mathbb{R}}\xi_{\textup{H}}(x),\ \xi_{8}(x)=\mathbb{R}^{4}\otimes_{\mathbb{R}}\xi_{\textup{H}}^{-}(x)\ \ x\in\partial_{\infty}\pi_{1}(S_{g}),$
are, by their definition, $\rho$-equivariant, continuous and transverse. Also for every $x\in\partial_{\infty}\pi_{1}(S_{g})$ we have $\xi_{1}(x)\in\xi_{4}(x)$, and hence $\xi_{1}$ is injective. It follows that $\xi_{1}(\partial_{\infty}\pi_{1}(S))=\Lambda_{\rho(\pi_{1}(S_{g}))}\cong S^{1}$. For $x\neq y$ the projective line segment $[\xi_{\textup{H}}(x),\xi_{\textup{H}}(y)]$ intersects $\Lambda_{\rho_{\textup{H}}(\Gamma)}$ at the set $\{\xi_{\textup{H}}(x),\xi_{\textup{H}}(y)\}$, hence $[\xi_{1}(x),\xi_{1}(y)]\cap\Lambda_{\rho(\Gamma)}=\{\xi_{1}(x),\xi_{1}(y)\}$.
The choice of the element $\gamma\in\pi_{1}(S_{g})$ shows that $\rho(\gamma)$ cannot be $P_{k}$-proximal for $k=2,4,6$, so $\rho$ is not $P_{k}$-Anosov for these values of $k$. Let $g\in\pi_{1}(S_{g})$ be a non-trivial element. The infinite sequence of elements $(\phi_{\ast}^{(n)}(g))_{n\in\mathbb{N}}$ has the property that $(|\phi_{\ast}^{(n)}(g)|_{\infty})_{n\in\mathbb{N}}$ is unbounded and there exists $M>0$ such that $\frac{\ell_{1}(\rho_{\textup{Fiber}}(\gamma_{n}))}{\ell_{2}(\rho_{\textup{Fiber}}(\gamma_{n}))}\leqslant M$ for every $n\in\mathbb{N}$. Then, it is easy to check that the ratio $\frac{\ell_{i}(\rho(\gamma_{n}))}{\ell_{i+1}(\rho(\gamma_{n}))}$ are uniformly bounded for $i=1,3,5$, so $\rho$ is not $P_{k}$-Anosov for $k=1,3,5$. ∎
@@
Necessity of the Cartan property. The representation $\rho\times\rho_{\textup{H}}:\pi_{1}(S_{g})\rightarrow\mathsf{SL}_{15}(\mathbb{R})$ (where $\rho$ and $\rho_{\textup{H}}$ are from Example 10.1) is $P_{1}$-divergent and admits a pair of continuous, equivariant, compatible and transverse maps $\xi^{+}:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathbb{P}(\mathbb{R}^{15})$ and $\xi^{-}:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathsf{Gr}_{14}(\mathbb{R}^{15})$ induced from the Anosov limit maps of $\rho_{\textup{H}}$. However, $\rho\times\rho_{\textup{H}}$ is not $P_{1}$-Anosov since $\rho$ cannot uniformly dominate $\rho_{\textup{H}}$. This shows that the assumption of the Cartan property for the map $\xi^{+}$ in Theorem 1.1 is necessary.

### v4

@@ 10. Examples and counterexamples
In this section, we discuss examples of representations of surface groups enjoying some of the properties of Anosov representations which are not $P_{1}$-Anosov. The examples show that the assumptions of the main results of this paper are necessary. Throughout this section, $g\in\mathbb{N}$ denotes the genus of a surface and $S_{g}$ denotes the (topological) closed orientable surface of genus $g\geq 2$.
Recall that for a subgroup $H$ of $\mathsf{GL}_{d}(\mathbb{R})$, containing a $P_{1}$-proximal element, we denote by $\Lambda_{H}$ its $P_{1}$-proximal limit set in $\mathbb{P}(\mathbb{R}^{d})$.
@@
  1. (1)
    $\rho$ is a quasi-isometric embedding, $P_{1}$-divergent and preserves a properly convex domain $\Omega$ of $\mathbb{P}(\mathbb{R}^{12})$.
  2. (2)
    $\rho$ admits continuous, injective, $\rho$-equivariant maps
    $(\xi_{1},\xi_{11}):\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathbb{P}(\mathbb{R}^{12})\times\mathsf{Gr}_{11}(\mathbb{R}^{12})$
    satisfying the Cartan property. The proximal limit set of $\rho(\pi_{1}(S_{g}))$ in $\mathbb{P}(\mathbb{R}^{12})$ is $\xi_{1}(\partial_{\infty}\pi_{1}(S_{g}))$ and does not contain projective line segments.
  3. (3)
    $\rho$ admits continuous, $\rho$-equivariant maps
    $(\xi_{4},\xi_{8}):\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathsf{Gr}_{4}(\mathbb{R}^{12})\times\mathsf{Gr}_{8}(\mathbb{R}^{12})$
    which are transverse.
  4. (4)
    $\rho$ is not $P_{k}$-Anosov for any $k=1,\ldots,11$.
The previous example shows that the assumption of transversality in Theorem 1.1 is necessary. Moreover, the maps $\xi_{4}$ and $\xi_{8}$ are transverse although $\rho$ is not $P_{4}$-Anosov, therefore the Zariski density assumption in Theorem 1.3 cannot be dropped.
@@
Let $g\geq 2$ and $\phi:S_{g}\rightarrow S_{g}$ a pseudo-Anosov homeomorphism. The mapping torus $M_{\phi}$ of $S_{g}$ with respect to $\phi$ is a closed $3$-manifold whose fundamental group is isomorphic to the HNN extension
$\pi_{1}(M_{\phi})=\Big\langle\pi_{1}(S_{g}),t\ \Big|\ tht^{-1}=\phi_{\ast}(h),\ h\in\pi_{1}(S_{g})\Big\rangle$
@@
where $\phi_{\ast}$ is a representative of the well-defined outer automorphism of $\pi_{1}(S_{g})$, induced by $\phi$. Thurston in [41] (see also Otal [39]) proved that there exists a convex cocompact representation $\rho_{0}:\pi_{1}(M_{\phi})\rightarrow\mathsf{PO}(3,1)$. The representation $\rho_{0}$ lifts to a $P_{1}$-Anosov representation in $\mathsf{SL}_{4}(\mathbb{R})$ which we continue to denote by $\rho_{0}$ and let $\rho_{\textup{Fiber}}:=\rho_{0}|_{\pi_{1}(S_{g})}$. By a result of Cannon-Thurston [15], there exists a continuous $\pi_{1}(S_{g})$-equivariant surjection $\theta:\partial_{\infty}\pi_{1}(S_{g})\twoheadrightarrow\partial_{\infty}\pi_{1}(M_{\phi})$. By precomposing $\theta$ with the Anosov limit map of $\rho_{0}$ in $\mathbb{P}(\mathbb{R}^{4})$, we obtain a $\rho_{\textup{Fiber}}$-equivariant continuous map $\xi_{\textup{Fiber}}:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathbb{P}(\mathbb{R}^{4})$.
Fix a pants decomposition of $S_{g}$ and let $\gamma_{0}\in\pi_{1}(S_{g})$ be an element representing a separating simple closed curve on this decomposition. We claim that there is a Zariski dense, Hitchin representation $\rho_{\textup{H}}:\pi_{1}(S_{g})\rightarrow\mathsf{SL}_{3}(\mathbb{R})$ with $\ell_{1}(\rho_{\textup{H}}(\gamma_{0}))=\lambda^{2}$, $\rho_{\textup{H}}(\gamma_{0})=\textup{diag}(\lambda^{2},1,\lambda^{-2})$ and $\lambda:=\ell_{1}(\rho_{\textup{Fiber}}(\gamma_{0}))$. To see this, using the fixed pants decomposition of $S_{g}$, we can fix a discrete faithful representation $j_{0}:\pi_{1}(S_{g})\rightarrow\mathsf{SL}_{2}(\mathbb{R})$ such that the modulus of the first eigenvalue of $j_{0}(\gamma_{0})$ is equal to $\lambda$. By composing $j_{0}$ with the irreducible representation $\textup{sym}^{2}:\mathsf{SL}_{2}(\mathbb{R})\rightarrow\mathsf{SL}_{3}(\mathbb{R})$, we obtain the Fuchsian representation $\textup{sym}^{2}j_{0}$ such that $\textup{sym}^{2}j_{0}(\gamma_{0})$ is conjugate to the matrix $\textup{diag}(\lambda^{2},1,\lambda^{-2})$. Then bending along the curve representing $\gamma_{0}$, gives a Zariski dense Hitchin representation $\rho_{\textup{H}}:\pi_{1}(S_{g})\rightarrow\mathsf{SL}_{3}(\mathbb{R})$, arbitrarily close to $\textup{sym}^{2}j_{0}$, with $\rho_{\textup{H}}(\gamma_{0})=\textup{sym}^{2}j_{0}(\gamma_{0})$.
@@
We claim that $\rho=\rho_{\textup{Fiber}}\otimes\rho_{\textup{H}}:\pi_{1}(S_{g})\rightarrow\mathsf{SL}_{12}(\mathbb{R})$ satisfies the required properties. Consider $\otimes:\mathsf{SO}(3,1)\times\mathsf{SL}_{3}(\mathbb{R})\rightarrow\mathsf{SL}_{12}(\mathbb{R})$ the irreducible tensor product representation $(g_{1},g_{2})\mapsto g_{1}\otimes g_{2}$. Let $G$ be the Zariski closure of $\rho_{\textup{Fiber}}\times\rho_{\textup{H}}$ into $\mathsf{SO}(3,1)\times\mathsf{SL}_{3}(\mathbb{R})$. Note that the projection of the identity component $G^{0}$ into $\mathsf{SO}(3,1)$ (resp. $\mathsf{SL}_{3}(\mathbb{R})$) is normalized by $\rho_{\textup{Fiber}}(\pi_{1}(S_{g}))$ (resp. $\rho_{\textup{H}}(\pi_{1}(S_{g}))$), so it has to be surjective. Since the Zariski closures of $\rho_{\textup{Fiber}}$ and $\rho_{\textup{H}}$ are simple and not locally isomorphic, it follows by Goursat’s lemma that $G=\mathsf{SO}(3,1)\times\mathsf{SL}_{3}(\mathbb{R})$. We conclude that $\rho$ is strongly irreducible.
@@
We obtain a properly convex domain $\Omega$ of $\mathbb{P}(\mathbb{R}^{12})$ preserved by $\rho(\pi_{1}(S_{g}))$ as follows. Let $\Omega_{1}$ and $\Omega_{2}$ be properly convex domains of $\mathbb{P}(\mathbb{R}^{4})$ and $\mathbb{P}(\mathbb{R}^{3})$ preserved by $\rho_{\textup{Fiber}}(\pi_{1}(S_{g}))$ and $\rho_{\textup{H}}(\pi_{1}(S_{g}))$ respectively, and $\Omega_{i}^{\prime}$ a properly convex cone lifting $\Omega_{i}$ for $i=1,2$. The compact set
$\mathcal{C}:=\Big\{[u_{1}\otimes u_{2}]\in\mathbb{P}(\mathbb{R}^{4}\otimes\mathbb{R}^{3}):u_{1}\in\overline{\Omega_{1}^{\prime}},u_{2}\in\overline{\Omega_{2}^{\prime}}\Big\}$
@@
is connected, spans $\mathbb{R}^{12}$ and is contained in an affine chart $\mathbb{A}\subset\mathbb{P}(\mathbb{R}^{4}\otimes\mathbb{R}^{3})\simeq\mathbb{P}(\mathbb{R}^{12})$. We finally take $\Omega$ to be the interior of the convex hull of $\mathcal{C}$ in $\mathbb{A}$.
The representations $\rho_{\textup{Fiber}}$ and $\rho_{\textup{H}}$ are $P_{1}$-divergent hence $\rho$ is also $P_{1}$-divergent as
$\sigma_{1}(\rho(\gamma))$ $=\sigma_{1}(\rho_{\textup{Fiber}}(\gamma))\sigma_{1}(\rho_{\textup{H}}(\gamma))\ \ \forall\gamma\in\Gamma$
$\sigma_{2}(\rho(\gamma))$ $=\max\big\{\sigma_{1}(\rho_{\textup{Fiber}}(\gamma))\sigma_{2}(\rho_{\textup{H}}(\gamma)),\sigma_{1}(\rho_{\textup{H}}(\gamma))\big\}\ \ \forall\gamma\in\Gamma.$
@@
In addition, since $\rho_{\textup{H}}$ is a quasi-isometric embedding, we deduce that $\rho$ is also a quasi-isometric embedding. Let $\xi_{\textup{H}}:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathbb{P}(\mathbb{R}^{3})$ and $\xi_{\textup{H}}^{-}:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathsf{Gr}_{2}(\mathbb{R}^{3})$ be the Anosov limit maps of $\rho_{\textup{H}}$. The map $\xi_{1}:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathbb{P}(\mathbb{R}^{12})$ defined as
$\xi_{1}(x)=\big[k_{x}e_{1}\otimes k_{x}^{\prime}e_{1}\big]$
@@
where $\xi_{\textup{Fiber}}(x)=[k_{x}e_{1}]$ and $\xi_{\textup{H}}(x)=[k_{x}^{\prime}e_{1}]$, is continuous and $\rho$-equivariant. Since $\rho$ is strongly irreducible, the proof of Corollary 4.6 shows that the map $\xi_{1}$ satisfies the Cartan property. The image of $\xi_{1}$ is the $P_{1}$-proximal limit set $\Lambda_{\rho(\pi_{1}(S_{g}))}$ of $\rho(\pi_{1}(S_{g}))$ in $\mathbb{P}(\mathbb{R}^{12})$, since $\Gamma$ acts minimally on $\partial_{\infty}\pi_{1}(S_{g})$. Similarly, the dual reprsentation $\rho^{\ast}=\rho_{\textup{Fiber}}^{\ast}\otimes\rho_{\textup{H}}^{\ast}$ admits a $\rho^{\ast}$-equivariant map $\xi_{1}^{\ast}:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathbb{P}(\mathbb{R}^{12})$, so we obtain the $\rho$-equivariant map $\xi_{11}$.
@@
$\xi_{4}(x)=\mathbb{R}^{4}\otimes\xi_{\textup{H}}(x),\ \xi_{8}(x)=\mathbb{R}^{4}\otimes\xi_{\textup{H}}^{-}(x)\ \ x\in\partial_{\infty}\pi_{1}(S_{g}),$
are, by their definition, $\rho$-equivariant, continuous and transverse. Also for every $x\in\partial_{\infty}\pi_{1}(S_{g})$ we have $\xi_{1}(x)\in\xi_{4}(x)$, hence $\xi_{1}$ is injective. It follows that $\xi_{1}(\partial_{\infty}\pi_{1}(S_{g}))=\Lambda_{\rho(\pi_{1}(S_{g}))}\cong S^{1}$. For $x\neq y$ the projective line segment $\xi_{\textup{H}}(x),\xi_{\textup{H}}(y)]$ intersects $\Lambda_{\rho_{\textup{H}}(\Gamma)}$ at the set $\{\xi_{\textup{H}}(x),\xi_{\textup{H}}(y)\}$, hence $\xi_{1}(x),\xi_{1}(y)]\cap\Lambda_{\rho(\Gamma)}=\{\xi_{1}(x),\xi_{1}(y)\}$. To see this, assume for $x_{1},x_{2},x_{3}\in\partial_{\infty}\pi_{1}(S_{g})$, $x_{2}\neq x_{3}$, and $\xi_{\textup{H}}(x_{i})=[u_{i}]$ and $\xi_{\textup{Fiber}}(x_{i})=[v_{i}]$. If $v_{1}\otimes u_{1}\in\mathbb{R}v_{2}\otimes u_{2}+\mathbb{R}v_{3}\otimes u_{3}$, then $e_{i}\otimes u_{1}\in\mathbb{R}e_{i}\otimes u_{2}+\mathbb{R}e_{i}\otimes u_{3}$, where $i\in\{1,\ldots,4\}$ is any index such that $\langle v_{1},e_{i}\rangle\neq 0$. This implies $\xi_{\textup{H}}(x_{1})\in\xi_{\textup{H}}(x_{2})\oplus\xi_{\textup{H}}(x_{3})$, hence $x_{1}=x_{2}$ or $x_{1}=x_{3}$.
The choice of the element $\gamma_{0}\in\pi_{1}(S_{g})$ such that $\ell_{1}(\rho_{\textup{H}}(\gamma_{0}))=\lambda^{2}$, $\lambda=\ell_{1}(\rho_{\textup{Fiber}}(\gamma_{0}))$, shows that the moduli of eigenvalues of $\rho(\gamma_{0})=\rho_{\textup{Fiber}}(\gamma_{0})\otimes\rho_{\textup{H}}(\gamma_{0})$ in non-increasing order are
$\lambda^{3},\lambda^{2},\lambda^{2},\lambda,\lambda,1,1,\lambda^{-1},\lambda^{-1},\lambda^{-2},\lambda^{-2},\lambda^{-3}.$
Thus, $\rho(\gamma_{0})$ is not $P_{k}$-proximal for $k=2,4,6$ and $\rho$ is not $P_{k}$-Anosov for $k=2,4,6$. Let $\delta\in\pi_{1}(S_{g})$ be a non-trivial element. Since $\phi$ is pseudo-Anosov, the infinite sequence of elements $(\phi_{\ast}^{(n)}(\delta))_{n\in\mathbb{N}}\subset\pi_{1}(S_{g})$ has the property that $(|\phi_{\ast}^{(n)}(\delta)|_{\infty})_{n\in\mathbb{N}}$ is unbounded (where $|\cdot|_{\infty}$ is the stable translation length with respect to a fixed word metric on $\pi_{1}(S_{g})$). By the definition of $\rho_{\textup{Fiber}}$, as $\rho_{\textup{Fiber}}(\phi_{\ast}^{(n)}(\delta))$ is conjugate to $\rho_{0}(\delta)$ for every $n$, there is $M>0$ such that $\frac{\ell_{1}(\rho_{\textup{Fiber}}(\phi_{\ast}^{(n)}(\delta)))}{\ell_{2}(\rho_{\textup{Fiber}}(\phi_{\ast}^{(n)}(\delta)))}\leqslant M$ for every $n\in\mathbb{N}$. Then, it is straightforward to check that the ratios $\Big(\frac{\ell_{i}(\rho(\phi_{\ast}^{(n)}(\delta)))}{\ell_{i+1}(\rho(\phi_{\ast}^{(n)}(\delta)))}\Big)_{n\in\mathbb{N}}$ are uniformly bounded for $i=1,3,5$, so $\rho$ is not $P_{k}$-Anosov for $k=1,3,5$. ∎
@@
Necessity of the Cartan property. The representation $\rho\times\rho_{\textup{H}}:\pi_{1}(S_{g})\rightarrow\mathsf{SL}_{15}(\mathbb{R})$ (where $\rho$ and $\rho_{\textup{H}}$ are from Example 10.1) is $P_{1}$-divergent since
$\frac{\sigma_{1}((\rho\times\rho_{\textup{H}})(\gamma))}{\sigma_{2}((\rho\times\rho_{\textup{H}})(\gamma))}=\frac{\sigma_{1}(\rho_{\textup{Fiber}}(\gamma))\sigma_{1}(\rho_{\textup{H}}(\gamma))}{\max\big\{\sigma_{1}(\rho_{\textup{Fiber}}(\gamma))\sigma_{2}(\rho_{\textup{H}}(\gamma)),\sigma_{1}(\rho_{\textup{H}}(\gamma))\big\}}\longrightarrow+\infty$
as $|\gamma|_{\pi_{1}(S_{g})}\rightarrow+\infty$.
In addition, the product $\rho\times\rho_{\textup{H}}$ admits a pair of continuous, $(\rho\times\rho_{\textup{H}})$-equivariant, compatible and transverse maps $\xi^{+}:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathbb{P}(\mathbb{R}^{15})$ and $\xi^{-}:\partial_{\infty}\pi_{1}(S_{g})\rightarrow\mathsf{Gr}_{14}(\mathbb{R}^{15})$, induced from the Anosov limit maps of $\rho_{\textup{H}}$, i.e. $\xi^{+}(x)=\{0\}\times\xi_{\textup{H}}(x)$ and $\xi^{-}(x)=\mathbb{R}^{12}\times\xi_{\textup{H}}^{-}(x)$, $x\in\partial_{\infty}\pi_{1}(S_{g})$. However, $\rho\times\rho_{\textup{H}}$ is not $P_{1}$-Anosov since $\rho$ cannot uniformly dominate $\rho_{\textup{H}}$. This shows that the assumption of the Cartan property for the map $\xi^{+}$ in Theorem 1.1 is necessary.